# UniTok: A Unified Tokenizer for Visual Generation and Understanding

### v2

@@ Abstract
Visual generative and understanding models typically rely on distinct tokenizers to process images, presenting a key challenge for unifying them within a single framework. Recent studies attempt to address this by connecting the training of VQVAE (for autoregressive generation) and CLIP (for understanding) to build a unified tokenizer. However, directly combining these training objectives has been observed to cause severe loss conflicts. In this paper, we show that reconstruction and semantic supervision do not inherently conflict. Instead, the underlying bottleneck stems from limited representational capacity of discrete token space. Building on these insights, we introduce UniTok, a unified tokenizer featuring a novel multi-codebook quantization mechanism that effectively scales up the vocabulary size and bottleneck dimension. In terms of final performance, UniTok sets a new record of 0.38 rFID and 78.6% zero-shot accuracy on ImageNet. Besides, UniTok can be seamlessly integrated into MLLMs to unlock native visual generation capability, without compromising the understanding performance. Additionally, we show that UniTok favors cfg-free generation, reducing gFID from 14.6 to 2.5 on ImageNet 256$\times$256 benchmark. GitHub: https://github.com/FoundationVision/UniTok.
@@ 1 Introduction
In this work, we aim to design a unified visual tokenizer to bridge the gap in multimodal generation and understanding. Intuitively, this can be achieved by integrating CLIP supervision into VQVAE training, resulting in a discrete tokenizer capturing both fine-grained details and high-level semantics. However, we empirically find this training recipe confronts severe convergence issues [62] and largely falls behind the CLIP baseline in multimodal understanding (Figure 1). While prior studies commonly attribute these challenges to conflicts between semantic and pixel-level feature learning [62, 38, 59], recent progress in visual generation suggests the opposite, showing semantic regularization could benefit tokenizers in reconstruction-oriented training [64, 4, 19]. Such disparity motivates us question: Do reconstruction and semantic losses truly conflict in tokenizer training?
To study the problem, we conduct a comprehensive ablation on the unified tokenizer training paradigm (Figure 3), which yields several intriguing findings: First, we show that removing reconstruction supervision, which leads to a vector-quantized CLIP model, does not improve understanding performance compared to the unified tokenizer. This observation indicates that the performance gap between unified and CLIP tokenizers mainly arises from vector quantization, rather than conflicts between learning objectives; Further analysis reveals that this gap is driven by two key factors: token factorization, which projects tokens into a lower-dimensional space for code index lookup [66], and discretization. These operations are essential for vector quantization but inevitably compromise the expressiveness of visual tokens. We thus argue that the primary bottleneck of unified tokenizers lies in the limited representational capacity of discrete token space.
@@
Building upon these techniques, we train a unified tokenizer called UniTok to bridge visual generation and understanding. Through extensive experiments, we demonstrate that UniTok achieves comparable or even better performance to domain-specific tokenizers: On ImageNet evaluation, UniTok records an impressive 0.38 reconstruction FID and 78.6% zero-shot accuracy at 256$\times$256 resolution; In building unified MLLMs, UniTok enables the MLLM with native visual generation capabilities while maintaining decent understanding performance. It outperforms the Liquid [60] baseline with a VQGAN tokenizer by 5.5% on VQAv2 [13], 9.2% on TextVQA [43], and 339 points on MME [65]; In addition, we demonstrate that semantic supervision leads to improved latent space structure for autoregressive generation, i.e., for class-conditional image generation on ImageNet 256$\times$256, UniTok significantly reduces generation FID without classifier-free guidance from 14.6 to 2.5 under the LlamaGen [44] framework, which aligns with recent findings in diffusion modeling [19, 4, 64].
@@ 3 Method
In this section, we introduce UniTok, a unified tokenizer well-suited for both visual generation and understanding tasks. We start with a unified training recipe that integrates reconstruction (VQVAE) and semantic (CLIP) supervisions (Section 3.1). However, we find that simply combining both training objectives leads to severe performance degradation, which can be mainly attributed to limited representational capacity of discrete tokens (Section 3.2). To this end, we propose multi-codebook quantization and attention projection to enhance the latent feature space and derive unified visual representations (Section 3.3). An overview of the framework is presented in Figure 2.
@@ 3.2 Quantization Bottleneck
Despite being augmented with CLIP supervision, we find that the unified tokenizer exhibits unsatisfactory performance in visual understanding tasks, significantly lagging behind the commonly used CLIP tokenizer. To figure out the underlying cause of this underperformance, we break down the key components involved in training a unified tokenizer, as illustrated in Figure 3. Starting with the CLIP baseline, we provide a step-by-step walk-through of all changes in following paragraphs.
@@
Reconstruction Supervision. Finally, we integrate reconstruction losses into the training process to build a unified tokenizer, as outlined in Section 3.1. Previous literature suggests that loss conflict between VQVAE and CLIP is a major cause of performance degradation in joint training [62]. We observe a similar phenomenon where joint training results in sub-optimal ImageNet zero-shot classification accuracy and reconstruction FID compared to specialized training. However, surprisingly, we find that this degradation has negligible impacts on downstream understanding performance. Moreover, the degradation in classification accuracy and reconstruction FID diminishes after we improve the quantization methods (detailed in the next section). Based on these observations, we speculate that the perceived loss conflict is only a superficial issue, and the primary cause of the underperformance lies in the limited representational capacity of discrete tokens.
@@ 3.3 UniTok
Discussions. MCQ shares a similar concept with residual quantization (RQ) [18] in using multiple codes to quantize a token, but differs fundamentally in design philosophy: RQ follows a coarse-to-fine quantization order, whereas MCQ adopts a divide-and-conquer strategy. This distinction gives MCQ unique advantages when operating in high-dimensional latent spaces, where codes tend to become increasingly sparse. For instance, with a latent dimension of 64-d, we observe that MCQ’s quantization loss is 15 to 45 times lower than that of RQ. This is because MCQ partitions the original latent space into multiple low-dimensional subspaces for quantization. Our ablation study in Table 8 further confirms the superiority of MCQ in unified tokenizer training.
Attention projection. Existing VQ methods usually employ linear or convolutional projection layers for token factorization. But as shown in Figure 3, this over-simplified design fails to preserve rich semantics when compressing the feature dimensions, leading to degraded understanding performance. To alleviate this problem, we suggest adapting the multi-head attention modules for factorization. Specifically, instead of concatenating features from multiple heads after the attention calculation, we replace the concatenation operation with average pooling to realize channel compression. Figure 5 provides a detailed illustration of the adaptation. Despite its simplicity, we find this design effectively strengthens the representational power of factorized tokens and stabilizes training.
@@ 4.1 Implementation Details
MLLM Setup. We instantiate a unified MLLM described in Section 3.4 with the Llama-2-7B base model [53]. Following Liquid, we first pretrain the model on a mix of multimodal data, which is composed of 10M language data from DCLM [22], 30M internal MidJourney-style synthetic data, and 30M re-captioned image-text pairs from COYO [33] and Laion [42]. Subsequently, we finetune the model on 1.5M text-to-image data and 1.5M multimodal instruction tuning data introduced in Mini-Gemini [23]. Specifically, the learning rate is set to 5e-5 in the pretraining stage and 2e-5 in the finetuning stage. For visual understanding evaluation, we report results on standard VQA benchmarks including VQAv2 [13], GQA [14], TextVQA [43], POPE [24], MME [65], and MM-Vet [71]. For visual generation evaluation, we report results on GenAI-Bench [26] and MJHQ-30K [20].
@@ 4.2 Tokenizer Comparison
Table 1: Comparison on ImageNet reconstruction FID and zero-shot classification accuracy. rFID is measured at 256$\times$256 resolution with 16×16$\times$16 × downsample ratio. $\dagger$ indicates model using pretrained CLIP weights for initialization. $\ast$ indicates model trained on OpenImages.
@@
We benchmark UniTok on ImageNet using two primary metrics: Fréchet Inception Distance (FID) to evaluate reconstruction quality, and top-1 zero-shot accuracy to assess image-text alignment. The results are presented in Table 1. To provide a fair comparison with tokenizers trained on small datasets, we also train a version of UniTok on OpenImages [17] solely with reconstruction supervision. It can be seen that UniTok excels in reconstruction quality compared to both unified and domain-specific tokenizers, recording an impressive 0.38 rFID on ImageNet with 16×16$\times$16 × downsampling ratio. As a discrete tokenizer, UniTok even surpasses the continuous VAE tokenizer from Stable Diffusion v2.1 [41], showcasing the superiority of the proposed multi-codebook quantization. For the perception performance, we observe that randomly initialized UniTok demonstrates suboptimal zero-shot classification accuracy. This is expected as current training schedule (i.e., one epoch on 1.28B samples) is insufficient for CLIP training to fully converge. It can be seen that initializing the model with pretrained CLIP weights largely alleviates the problem, boosting the zero-shot accuracy from 70.8% to 78.6%. In complement to quantitative results, we provide examples of reconstructed images in Figure 6.
@@ 4.3 Class-Conditional Image Generation
Recent studies on diffusion models indicate that injecting semantics into VAE training leads to a better-structured latent space, considerably enhancing guidance-free generation performance [64, 4, 19]. To evaluate whether UniTok possesses similar properties, we test it within the LlamaGen framework for class-conditional image generation. As shown in Table 2, UniTok reduces the FID by 12.11 compared to the VQGAN baseline in CFG-free generation, under the same generator setup. This implies that UniTok learns a more structured code distribution, benefiting autoregressive modeling.
Table 2: Class-conditional image generation results on ImageNet 256$\times$256. : VQGAN tokenizer from LlamaGen. : Images are generated at 384$\times$384 resolution and then resized to 256$\times$256 for evaluation. ‘Pre.’: precision; ‘Rec.’: recall; ‘CFG’: classifier-free-guidance.
@@ 4.4 Unified Understanding and Generation
Understanding Performance. We evaluate the understanding performance of UniTok on diverse VQA benchmarks in Table 3. Our unified MLLM showcases clear advantages when compared to other unified models that also utilize a discrete visual tokenizer. Specifically, UniTok significantly outperforms the Chameleon model, which relies on a traditional VQVAE tokenizer, by 7.2% higher accuracy on VQAv2. Additionally, it surpasses VILA-U, another model with a unified tokenizer, by 3.3% in accuracy on the TextVQA benchmark and by a notable margin of 112 points on the MME-Perception scores. Furthermore, we can see that UniTok largely narrows the performance gap with MLLMs that incorporate continuous visual tokenizers. These strong results confirm the candidacy of UniTok as a unified visual tokenizer for multimodal models.
@@
Method LLM Token Type Res. VQAv2 GQA TextVQA POPE MME MM-Vet
@@
Generation Performance. Table 4 presents the text-to-image generation performance of our unified MLLM on GenAI-Bench (advanced prompts), a challenging benchmark that measures the alignment of generated images with text prompts in complex dimensions such as counting, differentiation, comparison, and understanding logical relationships. In this demanding context, our model not only consistently outperforms other autoregressive unified models, but also achieves competitive performance against domain experts (diffusion models) trained on billions of images. The strong results underscore the superiority of our unified MLLM in complex text-to-image generation tasks. We also provide results on GenAI-Bench (basic prompts) in Table 9 in Appendix.

### v3

@@ Abstract
Visual generative and understanding models typically rely on distinct tokenizers to process images, presenting a key challenge for unifying them within a single framework. Recent studies attempt to address this by connecting the training of VQVAE (for autoregressive generation) and CLIP (for understanding) to build a unified tokenizer. However, directly combining these training objectives has been observed to cause severe loss conflicts. In this paper, we show that reconstruction and semantic supervision do not inherently conflict. Instead, the underlying bottleneck stems from limited representational capacity of discrete token space. Building on these insights, we introduce UniTok, a unified tokenizer featuring a novel multi-codebook quantization mechanism that effectively scales up the vocabulary size and bottleneck dimension. In terms of final performance, UniTok sets a new record of 0.38 rFID and 78.6% zero-shot accuracy on ImageNet. Besides, UniTok can be seamlessly integrated into MLLMs to unlock native visual generation capability, without compromising the understanding performance. Additionally, we show that UniTok favors cfg-free generation, reducing gFID from 14.6 to 2.5 on ImageNet $\times256$ benchmark. GitHub: https://github.com/FoundationVision/UniTok.
@@ 1 Introduction
In this work, we aim to design a unified visual tokenizer to bridge the gap in multimodal generation and understanding. Intuitively, this can be achieved by integrating CLIP supervision into VQVAE training, resulting in a discrete tokenizer capturing both fine-grained details and high-level semantics. However, we empirically find this training recipe confronts severe convergence issues [61] and largely falls behind the CLIP baseline in multimodal understanding (Figure˜1). While prior studies commonly attribute these challenges to conflicts between semantic and pixel-level feature learning [61, 37, 58], recent progress in visual generation suggests the opposite, showing semantic regularization could benefit tokenizers in reconstruction-oriented training [63, 4, 19]. Such disparity motivates us question: Do reconstruction and semantic losses truly conflict in tokenizer training?
To study the problem, we conduct a comprehensive ablation on the unified tokenizer training paradigm (Figure˜3), which yields several intriguing findings: First, we show that removing reconstruction supervision, which leads to a vector-quantized CLIP model, does not improve understanding performance compared to the unified tokenizer. This observation indicates that the performance gap between unified and CLIP tokenizers mainly arises from vector quantization, rather than conflicts between learning objectives; Further analysis reveals that this gap is driven by two key factors: token factorization, which projects tokens into a lower-dimensional space for code index lookup [65], and discretization. These operations are essential for vector quantization but inevitably compromise the expressiveness of visual tokens. We thus argue that the primary bottleneck of unified tokenizers lies in the limited representational capacity of discrete token space.
@@
Building upon these techniques, we train a unified tokenizer called UniTok to bridge visual generation and understanding. Through extensive experiments, we demonstrate that UniTok achieves comparable or even better performance to domain-specific tokenizers: On ImageNet evaluation, UniTok records an impressive 0.38 reconstruction FID and 78.6% zero-shot accuracy at $\times256$ resolution; In building unified MLLMs, UniTok enables the MLLM with native visual generation capabilities while maintaining decent understanding performance. It outperforms the Liquid [59] baseline with a VQGAN tokenizer by 5.5% on VQAv2 [13], 9.2% on TextVQA [42], and 339 points on MME [64]; In addition, we demonstrate that semantic supervision leads to improved latent space structure for autoregressive generation, i.e., for class-conditional image generation on ImageNet $\times256$, UniTok significantly reduces generation FID without classifier-free guidance from 14.6 to 2.5 under the LlamaGen [43] framework, which aligns with recent findings in diffusion modeling [19, 4, 63].
@@ 3 Method
In this section, we introduce UniTok, a unified tokenizer well-suited for both visual generation and understanding tasks. We start with a unified training recipe that integrates reconstruction (VQVAE) and semantic (CLIP) supervisions (Section˜3.1). However, we find that simply combining both training objectives leads to severe performance degradation, which can be mainly attributed to limited representational capacity of discrete tokens (Section˜3.2). To this end, we propose multi-codebook quantization and attention projection to enhance the latent feature space and derive unified visual representations (Section˜3.3). An overview of the framework is presented in Figure˜2.
@@ 3.2 Quantization Bottleneck
Despite being augmented with CLIP supervision, we find that the unified tokenizer exhibits unsatisfactory performance in visual understanding tasks, significantly lagging behind the commonly used CLIP tokenizer. To figure out the underlying cause of this underperformance, we break down the key components involved in training a unified tokenizer, as illustrated in Figure˜3. Starting with the CLIP baseline, we provide a step-by-step walk-through of all changes in following paragraphs.
@@
Reconstruction Supervision. Finally, we integrate reconstruction losses into the training process to build a unified tokenizer, as outlined in Section˜3.1. Previous literature suggests that loss conflict between VQVAE and CLIP is a major cause of performance degradation in joint training [61]. We observe a similar phenomenon where joint training results in sub-optimal ImageNet zero-shot classification accuracy and reconstruction FID compared to specialized training. However, surprisingly, we find that this degradation has negligible impacts on downstream understanding performance. Moreover, the degradation in classification accuracy and reconstruction FID diminishes after we improve the quantization methods (detailed in the next section). Based on these observations, we speculate that the perceived loss conflict is only a superficial issue, and the primary cause of the underperformance lies in the limited representational capacity of discrete tokens.
@@ 3.3 UniTok
Discussions. MCQ shares a similar concept with residual quantization (RQ) [18] in using multiple codes to quantize a token, but differs fundamentally in design philosophy: RQ follows a coarse-to-fine quantization order, whereas MCQ adopts a divide-and-conquer strategy. This distinction gives MCQ unique advantages when operating in high-dimensional latent spaces, where codes tend to become increasingly sparse. For instance, with a latent dimension of 64-d, we observe that MCQ’s quantization loss is 15 to 45 times lower than that of RQ. This is because MCQ partitions the original latent space into multiple low-dimensional subspaces for quantization. Our ablation study in Table˜8 further confirms the superiority of MCQ in unified tokenizer training.
Attention projection. Existing VQ methods usually employ linear or convolutional projection layers for token factorization. But as shown in Figure˜3, this over-simplified design fails to preserve rich semantics when compressing the feature dimensions, leading to degraded understanding performance. To alleviate this problem, we suggest adapting the multi-head attention modules for factorization. Specifically, instead of concatenating features from multiple heads after the attention calculation, we replace the concatenation operation with average pooling to realize channel compression. Figure˜6 provides a detailed illustration of the adaptation. Despite its simplicity, we find this design effectively strengthens the representational power of factorized tokens and stabilizes training.
@@ 4.1 Implementation Details
MLLM Setup. We instantiate a unified MLLM described in Section˜3.4 with the Llama-2-7B base model [52]. Following Liquid, we first pretrain the model on a mix of multimodal data, which is composed of 10M language data from DCLM [22], 30M internal MidJourney-style synthetic data, and 30M re-captioned image-text pairs from COYO [32] and Laion [41]. Subsequently, we finetune the model on 1.5M text-to-image data and 1.5M multimodal instruction tuning data introduced in Mini-Gemini [23]. Specifically, the learning rate is set to 5e-5 in the pretraining stage and 2e-5 in the finetuning stage. For visual understanding evaluation, we report results on standard VQA benchmarks including VQAv2 [13], GQA [14], TextVQA [42], POPE [24], MME [64], and MM-Vet [70]. For visual generation evaluation, we report results on GenAI-Bench [26] and MJHQ-30K [20].
@@ 4.2 Tokenizer Comparison
Table 1: Comparison on ImageNet reconstruction FID and zero-shot classification accuracy. rFID is measured at $\times256$ resolution with $16\times$ downsample ratio. $\dagger$ indicates model using pretrained CLIP weights for initialization. $\ast$ indicates model trained on OpenImages.
@@
We benchmark UniTok on ImageNet using two primary metrics: Fréchet Inception Distance (FID) to evaluate reconstruction quality, and top-1 zero-shot accuracy to assess image-text alignment. The results are presented in Table˜1. To provide a fair comparison with tokenizers trained on small datasets, we also train a version of UniTok on OpenImages [17] solely with reconstruction supervision. It can be seen that UniTok excels in reconstruction quality compared to both unified and domain-specific tokenizers, recording an impressive 0.38 rFID on ImageNet with $16\times$ downsampling ratio. As a discrete tokenizer, UniTok even surpasses the continuous VAE tokenizer from Stable Diffusion v2.1 [40], showcasing the superiority of the proposed multi-codebook quantization. For the perception performance, we observe that randomly initialized UniTok demonstrates suboptimal zero-shot classification accuracy. This is expected as current training schedule (i.e., one epoch on 1.28B samples) is insufficient for CLIP training to fully converge. It can be seen that initializing the model with pretrained CLIP weights largely alleviates the problem, boosting the zero-shot accuracy from 70.8% to 78.6%. In complement to quantitative results, we provide examples of reconstructed images in Figure˜4.
@@ 4.3 Class-Conditional Image Generation
Recent studies on diffusion models indicate that injecting semantics into VAE training leads to a better-structured latent space, considerably enhancing guidance-free generation performance [63, 4, 19]. To evaluate whether UniTok possesses similar properties, we test it within the LlamaGen framework for class-conditional image generation. As shown in Table˜2, UniTok reduces the FID by 12.11 compared to the VQGAN baseline in CFG-free generation, under the same generator setup. This implies that UniTok learns a more structured code distribution, benefiting autoregressive modeling.
Table 2: Class-conditional image generation results on ImageNet $\times256$. : VQGAN tokenizer from LlamaGen. : Images are generated at $\times384$ resolution and then resized to $\times256$ for evaluation. ‘Pre.’: precision; ‘Rec.’: recall; ‘CFG’: classifier-free-guidance.
@@ 4.4 Unified Understanding and Generation
Understanding Performance. We evaluate the understanding performance of UniTok on diverse VQA benchmarks in Table˜3. Our unified MLLM showcases clear advantages when compared to other unified models that also utilize a discrete visual tokenizer. Specifically, UniTok significantly outperforms the Chameleon model, which relies on a traditional VQVAE tokenizer, by 7.2% higher accuracy on VQAv2. Additionally, it surpasses VILA-U, another model with a unified tokenizer, by 3.3% in accuracy on the TextVQA benchmark and by a notable margin of 112 points on the MME-Perception scores. Furthermore, we can see that UniTok largely narrows the performance gap with MLLMs that incorporate continuous visual tokenizers. These strong results confirm the candidacy of UniTok as a unified visual tokenizer for multimodal models.
@@
Method LLM Token Type Res. VQAv2 GQA TextVQA POPE MME $M$-Vet
@@
Generation Performance. Table˜4 presents the text-to-image generation performance of our unified MLLM on the GenEval benchmark. We show that UniTok not only outperforms most of the unified MLLMs, but also demonstrates competitive performance against domain experts (diffusion models) trained on billions of images. Besides, UniTok achieves non-trivial improvements over Liquid while using exactly the same set of text-to-image training data, highlighting the importance of a unified tokenizer. We also provide results on GenAI-Bench in Table˜10 and Table˜11 in Appendix.
Table 4: Comparison with other visual generation methods on the GenEval benchmark.
[…]
SD v2.1 [39] Diffusion 2000M 0.98 0.51 0.44 0.85 0.07 0.17 0.50
SD-XL [36] Diffusion 2000M 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 [2] Diffusion 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Show-o [62] Discrete Diff. 36M 0.95 0.52 0.49 0.82 0.11 0.28 0.53
2.0B 0.98 0.80 0.66 0.84 0.31 0.50 0.68
LWM [27] Autoregressive 0.93 0.41 0.46 0.79 0.09 0.15 0.47
Janus [58] Autoregressive 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Liquid [59] Autoregressive 30M 0.98 0.73 0.32 0.76 0.17 0.37 0.55
UniTok Autoregressive 30M 0.99 0.71 0.36 0.79 0.26 0.45 0.59
We further evaluate the quality of images generated by our model on the MJHQ-30K benchmark, details of which are presented in Table˜5. Notably, as this benchmark primarily relies on the FID score
Table 5: Results on MJHQ-30K.
[…]
SD-XL [36] Diffusion 1024 9.55
PixArt [6] Diffusion 1024 6.14
Playground [20] Diffusion 1024 4.48
Liquid [59] Autoregressive 512 5.47
Janus [58] Autoregressive 384 10.10
LWM [27] Autoregressive 256 17.77
Show-o [62] Discrete Diff. 256 15.18
VILA-U [61] Autoregressive 256 12.81
UniTok Autoregressive 256 7.46
for evaluation, high-resolution images are preferred because they potentially capture more fine-grained details. Despite this makes FID across different resolutions less comparable, we show that our model achieves impressive performance even at the the smallest resolution, showcasing its ability to generate high-quality, detail-rich images.
We present some examples of the images generated by our model in Figure˜5, using text prompts sampled from MJHQ-30K. The visualization results demonstrate our model is capable of synthesizing photo-realistic and visually appealing images. Moreover, the model is able to comprehend a wide spectrum of concepts, such as ‘Vincent van Gogh painting style’ and ‘bitcoin’, and flexibly combine these concepts to synthesize creative images.
4.5 Ablation Studies
Impact of Supervision Types. To ablate the impact of contrastive and reconstruction losses in UniTok training, we conduct experiments on tokenizers trained with different supervision types, as shown in Table˜6. It is worth noting that all the tokenizers are vector-quantized even though some do not have reconstruction supervision. First, we show that reconstruction-oriented tokenizer significantly lags behind tokenizers with contrastive supervision in visual understanding performance. This observation evidences the limitations of traditional VQVAE. Second, we demonstrate that reconstruction and contrastive training objectives do not inherently conflict, or can be addressed by enhancing discrete feature space. With multi-codebook quantization, the jointly trained tokenizer not only exhibits understanding performance on par with the tokenizer trained solely with contrastive loss, but also slightly improves generation performance over the reconstruction-oriented tokenizer.
Supervision Generation Understanding
rFID $\downarrow$ gFID $\downarrow$ VQAv2 GQA SciQA TextVQA POPE MME
Contrastive 68.95 56.89 65.64 49.89 82.34 1373
Reconstruction 0.82 3.59 56.33 47.53 63.26 43.65 77.09 902
Recon. + Contra. 0.72 3.26 69.14 56.06 65.25 49.22 81.42 1333
MCQ v.s. RQ. Following discussions in Section˜3.3, we provide an apple-to-apple comparison between multi-codebook quantization and residual quantization in Table˜8. For fair comparisons, we directly implement RQ on our codebase, keeping all the training settings the same as UniTok. Both tokenizers are trained on a 512M subset of DataComp-1B. Notably, unlike MCQ, RQ by default uses a shared large codebook. To keep the global codebook size the same as UniTok, we set the codebook size of RQ to 32768. It can be seen that RQ demonstrates inferior reconstruction performance and lower classification accuracy compared to MCQ under the high bottleneck dimension (64-d) setting.
Number of Sub-Codebooks. To gain deeper insights into multi-codebook quantization, we evaluate how tokenizer performance changes with the number of sub-codebooks in Table˜8. Specifically, the size of a codebook is denoted as $A\times B$, where $A$ is the number of sub-codebook and $B$ is the size of sub-codebook. For rFID evaluation, we train the tokenizer solely with reconstruction loss on OpenImages [17], and evaluated it on ImageNet ($256\times 256$) validation set. While for ImageNet zero-shot accuracy evaluation, the tokenizer is trained on DataComp-1B 128m subset using only contrastive loss. Given a constant global codebook size, we see that increasing the number of sub-codebooks consistently improves reconstruction FID and classification accuracy. This indicates that MCQ generally benefits vector-quantized models, independent of the training objectives.
Table 7: MCQ v.s. RQ.
[…]
RQ $\times16$$\times$8 64 3.46 58.8
MCQ $\times16$$\times$8 64 0.55 63.7
Table 8: Ablation on number of sub-codebooks.
[…]
rFID $\downarrow$ 1.50 0.98 0.54 0.33
Accuracy $41.0\%$ $3.9\% 44$$4.7\%$ $46.1\%$
CLIP Weight Initialization. We notice that higher ImageNet accuracy does not guarantee superior downstream performance. In Table˜9, we ablate the impact of CLIP weight initialization on visual understanding performance. Specifically, we adopt the classic LLaVA framework for evaluation, replacing the original CLIP tokenizer with UniTok while keeping all other the training settings unchanged. One tokenizer is initialized with the pretrained ViTamin-L-256 [5] weights, while the other is randomly initialized. To our surprise, UniTok that is trained from scratch surpasses the one initialized with pretrained CLIP weights, despite the latter actually achieves better zero-shot classification accuracy. This suggests downstream VQA performance may not be highly correlated with ImageNet classification accuracy. More importantly, it also implies that CLIP weight initialization may serve as a negative prior for unified tokenizers, as the unified visual feature space could drastically differ from CLIP feature space.
Tokenizer VQAv2 GQA TextVQA POPE MME
UniTok 69.9 56.2 49.3 81.2 1331
UniTok 72.4 58.2 51.6 82.4 1392
5 Limitations and Conclusion
Acknowledgments
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Themebased Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). We sincerely thank Seng Ye for the insightful discussions and valuable contributions to this project.
Appendix A Attention Projection Modules
Figure˜6 illustrates the adaptations we made to the tradition MHA module, which enables channel compression and expansion in token factorization.
Appendix B More Generation Results
We provide the results on GenAI-Bench in Table˜10 and Table˜11. UniTok consistently delivers superior generation performance on this benchmark.
[…]
Spatial Action Part
SD v2.1 [39] Diffusion 2000M 0.80 0.79 0.76 0.77 0.80 0.78
SD-XL [36] Diffusion 2000M 0.84 0.84 0.82 0.83 0.89 0.83
Midjourney v6 Diffusion 0.88 0.87 0.87 0.87 0.91 0.87
DALL-E 3 [2] Diffusion 0.91 0.90 0.92 0.89 0.91 0.90
Show-o [62] Discrete Diff. 36M 0.72 0.72 0.70 0.70 0.75 0.70
LWM [27] Autoregressive 0.63 0.62 0.65 0.63 0.70 0.63
VILA-U [61] Autoregressive 15M 0.78 0.78 0.77 0.78 0.79 0.76
Liquid [59] Autoregressive 30M 0.84 0.86 0.81 0.83 0.91 0.83
UniTok Autoregressive 30M 0.85 0.87 0.86 0.86 0.89 0.85